A) Two of the SI base quantities are mass and time. State three other SI base quantities.

B) A sphere of radius r is moving at speed v through air of density p. The resistive force F acting on the sphere is given by the expression

F= Br^2 pv^k

Where B and k are constants without units

I) state the SI base units of F, p and v

II) Use base units to determine the value of k
1
Expert's answer
2018-04-18T08:08:07-0400
A)
Three other SI base quantities are time, electric current and thermodynamic themperature.
B)
I) F=kg*m/s2, ρ=kg/m3, v=m/s.
II) (kg*m)/s^2 =m^2*kg/m^3 *(m/s)^k,(kg*m)/s^2 =kg*m^(k-1)/s^k . It is obvious that k=2.